OpenStudy (anonymous):

The length of a rectangle is twice its width. If the perimeter of the rectangle is 54yd , find its area.

OpenStudy (callisto):

Let l be the length. length of a rectangle is twice its width => width = length /2 = l/2 Perimeter = 2 (length + width) = 2 (l + l/2) = 54 3l = 54 l = 18 Area = length x width = l (1/2) = 18 ( 18/2) = 162

OpenStudy (anonymous):

cool thats what I got, your awesome man(: